OpenStudy (anonymous):

Will medal and fan Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

can u help with 4 more

OpenStudy (anonymous):

Yea sure

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

12.4 + (–2 + 1.6) A. 11 B. 12 C. 15 D. 16

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thx

OpenStudy (welshfella):

@Xander22 - please guide the user to the answer - according the the Code of Coduct

OpenStudy (welshfella):

- dont just give the answer

OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (anonymous):

What do you think it is, we'll go from there

OpenStudy (anonymous):

@Xander22 You might want to consider giving @Destinyrules1234 an explanation to why those are the correct answers. & the answer to your last question is q is 3, because -(3+14)=-3+-14 -3 +-14=-17 -(3+14) the negative sign in front of the two numbers makes both numbers negative so -(3+14) is the same as -3+-14

OpenStudy (anonymous):

so is it 3?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

last one :D What value of x makes the equation true? –(9 + (–17)) = –9 + x

OpenStudy (anonymous):

Yes it is 3 which is C but do you understand why it is 3? Or do you want me to go over it again step by step ?

OpenStudy (anonymous):

i think it is 7 because it is opposite

OpenStudy (anonymous):

17*

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

hello????

OpenStudy (anonymous):

–(9 - 17) = –9 + x –(-8) = –9 + x 8 = –9 + x Add 9 8+9 = 9–9 + x 17 = x Check it –(9 + (–17)) = -(9-17) = -(-8) = +8 –9 + x ==> -9+17 = +8